OpenStudy (anonymous):

Find the x-intercept of the parabola with vertex (4,75) and y-intercept (0,27) (x 1 ,y 1 ),(x 2 ,y 2 )

OpenStudy (anonymous):

\(y\) intercept is what you get when you replace \(x\) by \(0\) if your parabola is \(y=ax^2+bx+c\) the the \(y\) intercept is \(c\) which you know is \(27\)

OpenStudy (anonymous):

since the vertex is \((4,75)\) you know it looks like \[y=a(x-4)^2+75\] and now we can solve for \(a\) by replacing \(x\) by \(0\) and setting the result equal to \(27\)

OpenStudy (anonymous):

you get \[27=a(0-4)^2+75\] \[27=16a+75\] \[-48=16a\] \[-4=a\]

OpenStudy (anonymous):

damn that is wrong sorry, should be \(-3=a\)

OpenStudy (anonymous):

therefore \[y=-3(x-4)^2+75\] and now we can set it equal to zero and solve

OpenStudy (anonymous):

you get \[-3(x-4)^2+75=0\] \[-3(x-3)^2=-75\] \[(x-3)^2=25\] \[x-3=\pm5\] \[x=8\] or \[x=-2\]

OpenStudy (anonymous):

and the \(x\) intercepts are therefore \((-2,0)\) and \((8,0)\)